OpenStudy (anonymous):

Find the value of c such that the point (2, 5) lies on the graph of f(x) = cx^2 + 4x − 2

OpenStudy (anonymous):

Alternative Form: x (c x+4) = f(x)+2 & Derivative Form: a/ax(Cx^2 + 4x - 2)= 2 c x +4